### v4

# A Hurewicz-type Theorem for the Dynamic Asymptotic Dimension with Applications to Coarse Geometry and Dynamics

Abstract:
We prove a Hurewicz-type theorem for the dynamic asymptotic dimension originally introduced by Guentner, Willett, and Yu. Calculations of (or simply upper bounds on) this dimension are known to have implications related to cohomology of group actions and the $K$-theory of their transformation group $C^{*}$-algebras. Moreover, these implications are relevant to the current classification program for $C^{*}$-algebras. As a corollary of our main theorem, we show the dynamic asymptotic dimension of actions by groups on profinite completions along sequential filtrations by normal subgroups is often subadditive over extensions of groups, which shows that many such actions by elementary amenable groups are finite dimensional. We combine this extension theorem with other novel results relating the dynamic asymptotic dimension of such actions to the asymptotic dimension of corresponding box spaces. This allows us to give upper bounds on the asymptotic dimension of many box spaces (including those of infinitely-many groups with exponential growth). For some of these examples, we can also find lower bounds by utilizing the theory of ends of groups.
1 Introduction
When considering any invariant thought to behave like dimension, a natural question is whether it is subadditive over morphisms. That is, given a morphism $f:X\to Y$ of two objects, is the dimension of $X$ at most the dimension of $Y$ plus the largest dimension of a ‘fibre’ of $f$ (speaking loosely)? An affirmative answer to such a question is sometimes called a Hurewicz-type theorem, and speaks to the naturality and dimension-theoretic nature of such an invariant. Classical results of this kind include the rank-nullity theorem (for the category of vector spaces and linear maps) and the Hurewicz mapping theorem [engelking, Theorem 1.12.4, page 136] (for the category of topological spaces and continuous maps). Special cases include whether a notion of dimension is subadditive over products or extensions of objects.
More recently, Hurewicz-type theorems were developed for Gromov’s asymptotic dimension in [asymptotichurewicztheorem] and later in [Brodskiy2006AHT]. This paper uses similar methods as in [Brodskiy2006AHT], together with some original ideas, to prove a Hurewicz-type theorem for the dynamic asymptotic dimension (DAD) of Guentner, Willett, and Yu. This dimension theory was introduced in [dasdimGWY] as a dynamical analogue of the asymptotic dimension. It is related to conditions used in work on the Farrell-Jones conjecture on manifold topology [bartels] [bartels_et_al], as well as to the homology of group actions [bonicke2021dynamic] and the nuclear dimension of transformation group $C^{*}$-algebras [dasdimGWY, Theorem 8.6].
We state the appropriate Hurewicz-type theorem for the dynamic asymptotic dimension below. We use the definition of DAD for free actions (see 2.5), in the statement, but do not assume the actions are free (more on this in 2.7).
Theorem A.
Suppose $\Gamma\curvearrowright X$ and $\Lambda\curvearrowright Y$ are actions on compact metric spaces by homeomorphisms, and that $f:X\to Y$ and $\alpha:\Gamma\times X\to\Lambda$ are continuous functions such that $f(\gamma\cdot x)=\alpha(\gamma,x)\cdot f(x)$ for all $x\in X$ and $\gamma\in\Gamma$. If $\text{DAD}(f):=\sup_{A\subset Y:\text{DAD}(A)=0}\text{DAD}_{free}(f^{-1}(A))$, it holds that
$\text{DAD}_{free}(\Gamma\curvearrowright X)\leq\text{DAD}(f)+\text{DAD}_{free}(\Lambda\curvearrowright Y).$
Apart from being a natural permanence property for the dynamic asymptotic dimension, this theorem is interesting because it provides a new tool for showing certain actions have finite DAD. While it is thought that an action by an amenable group with finite asymptotic dimension should have finite DAD, relatively little progress has been made aside from the case where the acting group is virtually nilpotent [szabo2017rokhlin, Theorem 7.4]. Finiteness of the dynamic asymptotic dimension is important to the current classification program for $C^{*}$-algebras because of how DAD relates to nuclear dimension and computations of $K$-theory via the Baum-Connes and HK conjectures [Guentner2016DynamicAD] [bonicke2021dynamic]. New examples of finite dimensional actions add to the class of algebras which are both classifiable by their $K$-theoretic and tracial data and which admit extensive technology for computing $K$-theory. DAD also plays a special role for those approaching classification from a dynamical perspective, as it applies to actions with finite stabilizers which are not essentially free, and therefore not covered by Kerr’s notion of almost finitness (see for instance [MR4066584] and [https://doi.org/10.48550/arxiv.2107.05273]). The introduction of [https://doi.org/10.48550/arxiv.2201.03409] gives a precise statement of the classification theorem and additional references.
We give two applications of this new Hurewicz-type theorem. First, we can show that the DAD is preserved when one restricts to a finite-index subgroup of the acting group, similar to how the asymptotic dimension of a finite index subgroup is the same as that of the larger group. This further demonstrates that the DAD is a coarse dimension theory for dynamics.
Theorem B.
Suppose $\Gamma\curvearrowright X$ is a free action on a compact metric space by homeomorphisms and $\Lambda<\Gamma$ is a finite index subgroup. Then $\text{DAD}(\Gamma\curvearrowright X)=\text{DAD}(\Lambda\curvearrowright X)$.
We can also show that the DAD of groups acting on their profinite completions (sometimes called odometer actions) behaves subadditively when one has an extension of groups, at least when the extension has certain properties:
Theorem C.
Suppose $1\to\Delta\to\Gamma\to\Lambda\to 1$ is an exact sequence of countable groups and that $\Gamma$ is residually finite. Suppose $(N_{i})$ is a countable collection of finite index normal subgroups of $\Gamma$ which separates $\Delta/(\cap_{i}(N_{i}\cap\Delta))$ from $\Gamma/(\cap_{i}N_{i})$ (see LABEL:separable_subgroup_definition). This collection induces collections of finite index normal subgroups and hence odometers of $\Delta$ and $\Lambda$. Then if $\text{DAD}_{free}(\Delta\curvearrowright\widehat{\Delta}_{(N_{i}\cap\Delta)})=m$ and $\text{DAD}_{free}(\Lambda\curvearrowright\widehat{\Lambda}_{(N_{i}/\Delta\cap N_{i})})=n$, $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})\leq m+n$.
This extension theorem can further be applied to coarse geometry, specifically the study of box spaces of groups. These are spaces constructed out of finite quotients of a group. Just as the asymptotic dimension of a group itself can have interesting group-theoretic applications (see for instance [MR2431020]), there are likewise many correspondences between geometric properties of box spaces and group-theoretic properties. For example, coarse equivalence of box spaces encodes virtual isomorphism of groups [MR4091056], and amenability of groups is equivalent to property A of their box spaces [Roe2003LecturesOC, Propsotion 11.39]. Similar results also exist for a-T-menablity [https://doi.org/10.1112/blms/bdt045].
Box spaces provided some of the first examples of spaces which do not coarsely embed in hilbert space [Roe2003LecturesOC, Remark 11.27] and it has been shown that a group with a finite dimensional box space must be amenable [Roe2003LecturesOC, Remark 11.38 and Proposition 11.39]. Similar to how actions by finite dimensional amenable groups are thought to be finite dimensional, it is thought that a partial converse to this latter fact should also hold; that is, an amenable group with finite asymptotic dimension ought to have finite dimensional box spaces. Notwithstanding, there has been similarly little progress on this aside from the case of virtually nilpotent groups [boxspacesDT].
As it turns out, questions about the asymptotic dimension of box spaces are essentially special cases of questions about the dynamic asymptotic dimension, echoing a relationship between DAD and the asymptotic dimension of warped cones described in [warpedcones, Section 8]. More precisely, we prove the following in 3.
Theorem D.
Let $\Gamma$ be a countable group and $(N_{i})$ a sequence of finite index subgroups of $\Gamma$. Suppose $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$ is the action by left multiplication of $\Gamma$ on its profinite completion along $(N_{i})$. Then $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\text{asdim}\Box_{(N_{i})}\Gamma$.
This relationship allows us to prove a special case of an extension theorem for the asymptotic dimension of box spaces. In fact, such an extension theorem was our original motivation for investigating a Hurewicz-type theorem for the DAD. A proof of a similar result was attempted in [finnsell2015asymptotic], but the proof therein contains a gap which we discuss later in this section.
Theorem E.
Suppose $1\to\Delta\to\Gamma\xrightarrow{\pi}\Lambda\to 1$ is an exact sequence of countable groups and $(N_{i})$ is a sequence of finite index normal subgroups of $\Gamma$ which separates $\Delta/(\cap_{i}(N_{i}\cap\Delta))$ from $\Gamma/(\cap_{i}N_{i})$. This sequence induces sequences and hence box spaces of $\Delta$ and $\Lambda$. Then if $\text{asdim}\Box_{(N_{i}\cap\Delta)}\Delta=m$ and $\text{asdim}\Box_{(\pi(N_{i}))}\Lambda=n$, $\text{asdim}\Box_{(N_{i})}\Gamma\leq m+n$.
We have therefore established a special case of the (unproved) statement [finnsell2015asymptotic, Theorem A]. Our extra assumptions are that $\Gamma$ is countable, that the $N_{i}$ are normal, and that $(N_{i})$ is a sequence separating $\Delta$ from $\Gamma$. In [finnsell2015asymptotic, Theorem A], the family of subgroups are only assumed to be semiconjugacy separating which is weaker than assuming $(N_{i})$ to be a filtration by normal subgroups. We expect analogous techniques would work to prove these results for uncountable groups, but require translating everything to the language of coarse structures. It is worth remarking that a countable group may have uncountably-many finite index subgroups.
Something should be said about how our approach differs from that of [finnsell2015asymptotic]. We employ a similar general strategy in attempting to apply a Hurewicz-type theorem, and as our notation will indicate, we are using similar arguments to handle DAD as are used in [Brodskiy2006AHT] to handle asymptotic dimension. We also still need to overcome an analogous problem as in [finnsell2015asymptotic, Lemma 4.1(5)]. An example first provided by Jiawen Zhang and restated below shows the lemma just cited is false. We are very grateful to Jianchao Wu for helping to communicate this example and explain the problem that it highlights. There are even ‘worse’ counterexamples one could construct, but this one is very simple.
Example.
Let $6\mathbb{Z}$ and $9k\mathbb{Z}$ be subgroups of $\mathbb{Z}$ for $k$ odd. Let $6\mathbb{Z}/(6\mathbb{Z}\cap 9k\mathbb{Z})\to\mathbb{Z}/9k\mathbb{Z}$ be the natural inclusion (see the diagram in [finnsell2015asymptotic, Remark 4.2]). When considered with the quotient metrics induced from a metric on $\mathbb{Z}$, this inclusion distorts the length of elements by an arbitrarily large amount for $k$ sufficiently large.
In order to avoid this problem, we assume a kind of compatibility between the family $(N_{i})$ and the particular extension. More precisely, we assume $(N_{i})$ separates $\Delta$ from $\Gamma$ (see LABEL:separable_subgroup_definition for a definition). We will see in LABEL:baumslag-solitar_groups_section that there are still significant applications, as all elementary amenable groups are built up from extensions which are compatible with sufficiently large collections. In fact, for our motivating example of the Baumslag-Solitar groups $\text{BS}(1,n)$, we will see that any sequence of finite index subgroups satisfies the hypotheses of our extension theorem. We show that box spaces of $BS(1,n)$ along filtrations have asymptotic dimension $2$ for all $n$, which in particular exhibits an infinite class of groups with exponential growth but finite-dimensional box spaces. In obtaining the lower bound for this result, we compute the asymptotic dimension of $BS(1,n)$, which by itself involves some interesting methods in geometric group theory. Specifically the notion of ends of groups allows one to prove that a torsion free group of asymptotic dimension $1$ is free (as we will see, this is an immediate consequence of results already in the literature). We also deal with some examples of more complicated groups which have similar presentations or are in some sense ‘generalized Baumslag-Solitar groups’.
More generally, it is natural to try to use our extension theorem along similar lines as in [finnsell2015asymptotic, Section 4], which attempts to relate the asymptotic dimension of box spaces of residually finite, elementary amenable groups to their Hirsch length. While the author encountered difficulty implementing this idea because of the additional assumptions in the extension theorem, we will still discuss how the content of this paper could be used to recover similar results. Specifically, one can probably use the Hurewicz theorem for the asymptotic dimension together with the ideas in section LABEL:applications to obtain an extension theorem for box spaces which doesn’t require the collection $(N_{i})$ to be a sequence; then use this to prove some special cases of the claims found in [finnsell2015asymptotic, Section 4].
2 (Dynamic) Asymptotic Dimension
The following concept is originally due to Gromov, specifically [Gromov1991AsymptoticIO, Section 1.E]. There are many definitions, but we will use one throughout which makes the analogy with dynamic asymptotic dimension obvious:
Definition 2.1.
Let $(X,d)$ be a metric space. If $U\subset X$ is a subset, an $r$-chain in $U$ is a sequence $x_{0},\ldots,x_{n}$ of points in $U$ such that $d(x_{i},x_{i+1})\leq r$ for all $i$; and two points in $U$ are in the same $r$-component of $U$ if they are connected by an $r$-chain with all points contained in $U$. The subset $U$ is said to be $r$-connected if any two points in $U$ are connected by an $r$-chain in $U$. Finally, $\text{asdim}X\leq n$ means for all $r>0$ there is $M>0$ and a cover $\mathcal{U}=\{U_{0},\ldots,U_{n}\}$ of $X$ such that each $r$-component of each $U_{i}$ has diameter at most $M$. Such a cover is called a $(d,R,M)$-cover of $X$. The asymptotic dimension of $X$ is then the smallest integer $d$ such that $\text{asdim}X\leq d$, and is infinite if there is no such $d$.
Examples 2.2.
Many of the coarse spaces we consider will be graphs, considered as metric spaces by equipping their vertex set with the $l^{1}$-path metric. Finitely generated groups can be considered as metric spaces through the Cayley graph construction. More precisely, if $F\subset\Gamma$ is a finite generating set we denote by $C_{F}(\Gamma)$ the graph whose vertices are the elements of $\Gamma$ and which has an edge between $\gamma$ and $\delta$ if there is $f\in F$ such that $f\gamma=\delta$. For our purposes, we can always assume $F=F^{-1}$ so that we can think of $C_{F}(\Gamma)$ as an undirected graph. The $l^{1}$-path metric on the set of vertices is then the same as the word metric $d_{F}(\gamma,\delta)=|\gamma\delta^{-1}|_{F}$ (where $|\cdot|_{F}$ denotes the minimal length according to the generating set $F$). More generally, any countable group, if $\Gamma$ is a countable group, it has a right-invariant proper metric unique up to coarse equivalence [notes_on_prop_A, Lemma 2.3.3]. We will fix, from now on, some such metric on $\Gamma$ and note that the uniqueness up to coarse equivalence implies we can sensibly refer to $\text{asdim}\Gamma$.
Notation 2.3.
Throughout, $\Gamma\curvearrowright X$ and $\Lambda\curvearrowright Y$ will be actions on compact metric spaces by homeomorphisms.
We now define the dynamic asymptotic dimension of a group action, originally introduced by Guentner, Willett, and Yu in [dasdimGWY].
Definition 2.4.
Let $\Gamma\curvearrowright X$ be an action on a compact Hausdorff space by homeomorphisms. The dynamic asymptotic dimension $\Gamma\curvearrowright X$, denoted $\text{DAD}(\Gamma\curvearrowright X)$, is the smallest integer $d$ such that for all finite $F\subset\Gamma$ there exists an open cover $\mathcal{U}=\{U_{0},\ldots,U_{d}\}$ of $X$ such that, for all $0\leq j\leq d$, the set
$\Bigg{\{}\begin{array}[]{l|l}g\in\Gamma&g=f_{k}\cdots f_{1}\text{ and }\exists x\in U_{j}\\ &\text{ such that }f_{k_{0}}\cdots f_{1}\cdot x\in U_{j}\text{ }\forall\text{ }1\leq k_{0}\leq k\end{array}\Bigg{\}}$
is finite.
If $\Gamma\curvearrowright X$ is free, this definition is equivalent to the following:
Definition 2.5.
If $\Gamma\curvearrowright X$ is an action on a compact space by homeomorphisms, $\text{DAD}_{free}(\Gamma\curvearrowright X)$ is the smallest integer $d$ such that for all finite $F\subset\Gamma$ there is an open cover $\mathcal{U}=U_{0},\ldots,U_{d}$ such that for all $0\leq j\leq d$ and all $x\in U_{j}$, the set
[…]
is uniformly finite with respect to $x$ (i.e. the same upper bound applies for all choices of $x$).
The following terminology for DAD makes the analogy with asymptotic dimension clear, and will be critical to the proof of our main theorem. It is inspired by [Brodskiy2006AHT].
Definition 2.6.
Let $\mathcal{P}_{\text{fs}}(\Gamma)$ denote the collection of all finite, symmetric subsets of $\Gamma$ containing the identity. Let $X$ be a compact metrizable space, $\Gamma\curvearrowright X$ an action by homeomorphisms, $A\subset X$ a subset, and $F\in\mathcal{P}_{\text{fs}}(\Gamma)$. An $F$-chain in $A$ is a finite sequence $x_{0},\ldots,x_{n}$ of points in $A$ such that $x_{i+1}=f_{i+1}\cdot x_{i}$ with $f_{i+1}\in F$ ($0\leq i<n$). Two points $x,y\in A$ are in the same $F$-component of $A$ if they are connected by an $F$-chain in $A$ (this is an equivalence relation since $F$ is symmetric). We say $A$ is $F$-connected if any two points in $A$ can be connected by an $F$-chain. If $S\in\mathcal{P}_{\text{fs}}(\Gamma)$, we say $A\subset X$ is $S$-bounded if it is a subset of $S\cdot x$ for some $x\in X$. A cover of $A$ by $m+1$ sets which are open (respectively, closed) in the subspace topology on $A$ and whose $F$-components are $S$-bounded is called an open (respectively, closed) $(m,F,S)$-cover for $A$. Notice that in this case $S$ can always be taken to be $F^{r}$ for some $r>0$. In case we need to distinguish between different actions on $X$ (as we will in section LABEL:applications), we will refer to an (open/closed) $(m,F,S)$-cover of $A$ with respect to $\Gamma\curvearrowright X$.
An $n$-dimensional control function for $A$ (with respect to $\Gamma\curvearrowright X$) is a function $\mathcal{D}_{A}:\mathcal{P}_{\text{fs}}(\Gamma)\to\mathcal{P}_{\text{fs}}(\Gamma)$ such that for every $F\in\mathcal{P}_{\text{fs}}(\Gamma)$, there is an $(n,F,\mathcal{D}_{X}(F))$-cover for $A$ (again, with respect to $\Gamma\curvearrowright X$ if the action is ambiguous). Note that $\text{DAD}_{free}(\Gamma\curvearrowright X)\leq n$ iff there is an $n$-dimensional control function for $X$. We similarly define $\text{DAD}(A)$ to be the least integer $n\geq 0$ such that there exists an $n$-dimensional control function for $A$.
Remark 2.7.
Although the terminology in 2.6 relates specifically to the definition of DAD for free actions, we will use it even for actions which are not free. The Hurewicz-type theorem that we will prove in section 4 will be stated using the definition for free actions. Note, however, that by the definition for non-free actions, any action with infinite stabilizers will be infinite dimensional. Since the two definitions are equivalent for an action with uniformly-finite stabilizers, and actions with uniformly-finite stabilizers can easily be related to free actions by ignoring some torsion elements, we do not really lose anything by restricting our attention in this way. From now on we will only use the definition for free actions aside from a brief mention of the general definition in 3.
For actions on totally disconnected spaces, we can equivalently require the covers to be clopen, as shown below.
Lemma 2.8.
Suppose $\mathcal{V}=\{V_{0},\ldots,V_{d}\}$ is an open cover of a totally disconnected, compact space $X$. Then there is a clopen cover $\mathcal{W}=\{W_{0},\ldots,W_{d}\}$ such that $W_{i}\subset V_{i}$ for all $0\leq i\leq d$. In the context of the previous definition, this implies if $\mathcal{V}$ is a $(d,F,S)$-cover, then $\mathcal{W}$ is a $(d,F,S)$-cover.
Proof.
Given $\mathcal{V}$, there is a cover $\mathcal{U}=\{U_{i}\}_{i=0}^{d}$ by open sets such that $\overline{U_{i}}\subset V_{i}$ for $0\leq i\leq d$ (a more general statement is shown in [munkres], Lemma 41.6). Then since $X$ has a basis of clopen sets, we can find for each $i$ and each $x\in\overline{U}_{i}$ a clopen set $W_{i,x}$ containing $x$ and contained in $V_{i}$. For each $i$, the $W_{i,x}$’s are a cover of $\overline{U_{i}}$ (which is compact since $X$ is compact); and so there is a finite subcover, the union over which we denote $W_{i}$. This is a finite union of clopen sets and therefore clopen. Moreover, the collection $\{W_{i}\}_{i=0}^{d}$ is a cover of $X$, and $W_{i}\subset V_{i}$ for all $i$. ∎
Lemma 2.9.
Suppose $\Gamma\curvearrowright X$ is a free action. Then $\text{DAD}(\Gamma\curvearrowright X)\geq\text{asdim}\Gamma$. In fact, this inequality holds even if one modifies the definition of DAD to allow covers by Borel sets, or even arbitrary sets.
Proof.
Fix $r>0$. Let $F=B_{e}^{r}(\Gamma)$. Suppose $\text{DAD}(\Gamma\curvearrowright X)\leq d$. Then for some $S\in\mathcal{P}_{\text{fs}}(\Gamma)$, there is a $(d,F,S)$-cover, $\mathcal{V}$, for $\Gamma\curvearrowright X$. Let $\mathcal{O}(x)$ be the orbit of some $x\in X$. Identify $\mathcal{O}(x)$ with $\Gamma$ by $\gamma\cdot x\mapsto\gamma$. Then $\mathcal{V}$ gives rise to a cover $\mathcal{V}^{\prime}$ of $\Gamma$. Moreover, two points $y,y^{\prime}\in\mathcal{O}(x)$ are connected by an $F$ chain iff their corresponding elements in $\Gamma$ are connected by an $r$-chain. Therefore, the $r$-components of each $V\in\mathcal{V}^{\prime}$ are $S$-bounded, so $\mathcal{V}^{\prime}$ is a $(d,r,\text{diam}(S))$-cover for $C_{F}(\Gamma)$. ∎
Definition 2.10.
If $\{G_{n}\}_{n\in\mathbb{N}}$ is a countable collection of finite metric spaces, the coarse disjoint union $\sqcup_{n}G_{n}$ is the metric space with underlying set the disjoint union of the $G_{n}$ and metric $d(x,y):=d_{G_{n}}(x,y)$ if $x,y\in G_{n}$ and $d(x,y):=\text{diam}(G_{n})+\text{diam}(G_{m})$ if $x\in G_{n}$ and $y\in G_{m}$ with $m\neq n$. If $\Gamma$ is a group and $\{N_{i}\}$ is a collection of finite-index subgroups, we define $\Box_{(N_{i})}\Gamma:=\sqcup_{i}\Gamma/N_{i}$ where each $\Gamma/N_{i}$ has the metric induced from the right-invariant proper metric on $\Gamma$ we fixed earlier (i.e. the distance between cosets is the infimum over the distances between their different elements according to the original metric on $\Gamma$). This is called a box space of $\Gamma$. If the metric on $\Gamma$ is the word metric coming from some generating set and if the $N_{i}$ are normal, the induced metrics on quotients are the word metrics coming from the image of that generating set in the quotient.
Definition 2.11.
A filtration of $\Gamma$ is a collection $\{N_{i}\}$ of finite index subgroups of $\Gamma$ such that $\bigcap_{i}N_{i}=\{e\}$, the trivial subgroup. If $\Gamma$ is residually finite, the collection of all finite, index normal subgroups is then a filtration. Note that a countable, residually finite group has a filtration which is also a sequence, that is $N_{i}\supset N_{i+1}$.
Remark 2.12.
Our definition of coarse disjoint union is slightly different from the one found in [boxspacesDT, Section 2], and so the box space we associate to $\{N_{i}\}$ is not necessarily a coarse disjoint union over $\Gamma/N_{i}$ in their sense (as the space between different components may not tend to infinity). However, our definition of box space is the same as theirs in the case when $\Gamma$ is finitely generated and residually finite and $\{N_{i}\}$ is a filtration of $\Gamma$. Note also that we do not always require the subgroups $N_{i}$ to be normal, as this is not necessary for some of our results. We will at various times need to assume that the $N_{i}$ are normal and/or that they form a sequence, meaning $N_{i+1}\subset N_{i}$. We are therefore careful to repeat our assumptions in each statement.
Definition 2.13.
If $(G_{i})$ is a sequence of metric spaces, we say $\text{asdim}(G_{i})\leq d$ uniformly if for all $r>0$ there is an $R>0$ such that each $G_{i}$ admits a $(d,r,R)$-cover. We similarly say for a sequence $(\Gamma\curvearrowright X_{i})$ of actions on spaces that $\text{DAD}(\Gamma\curvearrowright X_{i})\leq d$ uniformly if for all $F\in\mathcal{P}_{\text{fs}}(\Gamma)$ there is $S\in\mathcal{P}_{\text{fs}}(\Gamma)$ such that there is a $(d,F,S)$-cover for each action. We can similarly define what it means for a general collection of metric spaces or actions to have asdim or DAD at most $d$ uniformly. In terms of control functions, this means having one control function which ‘works’ for all spaces or actions in the collection.
Lemma 2.14.
$\text{asdim}\Box_{(N_{i})}\Gamma\leq d$ if and only if $\text{asdim}(\Gamma/N_{i})\leq d$ uniformly.
Proof.
For one inequality, notice that a $(d,r,R)$-cover for $\Box_{(N_{i})}\Gamma$ gives rise to a $(d,r,R)$-cover for each $\Gamma/N_{i}$ (just take the intersection).
For the other, let $r>0$ and find a $(d,r,R)$-cover $\mathcal{U}^{i}=\{U^{i}_{0},\ldots,U^{i}_{d}\}$ of each $\Gamma/N_{i}$. Let $I_{r}=\{i\mid\text{diam}(\Gamma/N_{i})\leq r\}$. Then $\cup_{i\in I_{r}}\Gamma/N_{i}\subset\Box_{(N_{i})}\Gamma$ has diameter at most $2r$ (hence is 0-dimensional) and the family $\{\Gamma/N_{i}\subset\Box_{(N_{i})}\Gamma|i\notin I_{r}\}$ is $r$-disjoint. The desired inequality follows from [MR1808331, Theorem 1] (we can simply put together the separate covers of the $\Gamma/N_{i}$ for $i\notin I_{r}$). ∎
The following lemma shows that the dimension of groups and box spaces behaves nicely with respect to direct limits. While we will not really use it, it is relevant to the discussion of elementary-amenable groups in LABEL:baumslag-solitar_groups_section and moreover does not seem to be written down anywhere.
Lemma 2.15.
Suppose $G=\bigcup_{i}G_{i}$ is an increasing union of groups and $X=\Box_{(H_{j})}G=\bigcup_{i}X_{i}$ is an increasing union of box spaces where $X_{i}=\Box_{(H_{j}\cap G_{i})}G_{i}$ is a box space of $G_{i}$ (here the metrics on $X$ and $G_{i}$ are induced from the right invariant, proper metric on $G$, and the metric on $X_{i}$ is constructed using the metric on $G_{i}$). Then $\text{asdim}G=\sup_{i}\text{asdim}G_{i}$ and $\text{asdim}X=\sup_{i}\text{asdim}X_{i}$.
Proof.
Fix $R>0$. Then there is $N$ such that $B_{e}^{R}(G)\subset G_{N}$. Notice that $d_{G}(G_{N}\cdot g,G_{N}\cdot h)$ is either 0 or >R>R. This follows since if $d_{G}(G_{N}\cdot g,G_{N}\cdot h)\leq R$ then the two orbits have a nontrivial intersection and therefore coincide. Since different $G_{N}$-orbits are >R>R apart, we can cover each one separately using $\text{asdim}G_{N}+1$ sets. For the second claim, find $N$ as before. Then if $x,y\in G/H_{j}$ are in separate $G_{N}$-orbits, their representatives in $G$ must be in separate $G_{N}$ orbits, and so $d_{G/{H_{j}}}(x,y)>R$. We can therefore similarly cover each orbit separately using $\text{asdim}\Box_{(H_{j}\cap G_{N})}G_{N}+1$ sets. ∎
In addition to 2.9, we see that when the space being acted on is discrete, there is a very concrete relationship between the dynamic asymptotic dimension of the action and the asymptotic dimension of the orbits.
Lemma 2.16.
Suppose $\Gamma$ is a countable group with subgroup $N<\Gamma$ and recall we have equipped $\Gamma$ with a right-invariant, proper metric (see 2.2). Suppose further that $\Gamma/N$ is a quotient of $\Gamma$ with the metric induced from $\Gamma$. Consider the action $\Gamma\curvearrowright\Gamma/N$ on left cosets by left multiplication (this is an action even if $N$ is not normal). Then an $S$-bounded subset has diameter at most $\text{diam}(S)$ in $\Gamma$. Moreover, a subset $A\subset\Gamma/N$ with diameter at most $R$ is $B_{e}^{R}(\Gamma)$-bounded (recall that quotients do not increase the growth of a group). In particular, this implies a $(d,B_{e}^{r}(\Gamma),S)$-cover for $\Gamma\curvearrowright\Gamma/N$ is a $(d,r,\text{diam}(S))$-cover of $\Gamma/N$; and a $(d,r,R)$-cover of $\Gamma/N$ is a $(d,B_{e}^{r}(\Gamma),B_{e}^{R}(\Gamma))$-cover for $\Gamma\curvearrowright\Gamma/N$.
Proof.
Observe that $x,y\in A\subset\Gamma$ are connected by a $B_{e}^{r}(\Gamma)$-chain in A iff they are connected by an $r$-chain in $A\subset\Gamma$. This follows since $d_{\Gamma/N}(\gamma\cdot x,x)\leq r$ for any $\gamma\in B_{e}^{r}(\Gamma)$ and conversely any elements $\leq r$ away from each other in $\Gamma/N$ are connected by some $\gamma\in B_{e}^{r}(\Gamma)$. ∎
There is a familiar dynamical object which is reminiscent of a box space called an odometer. We will see in 3 that the asymptotic dimension of box spaces is closely related to the dynamic asymptotic dimension of odometers.
Definition 2.17.
Let $\Gamma$ be a countable group and $(N_{i})$ ($i\in I$) a countably infinite collection of finite-index subgroups directed by inclusion. The profinite completion of $\Gamma$ along $(N_{i})$ is the inverse limit ${\lim_{\leftarrow i}\Gamma/N_{i}}$ and we denote it $\widehat{\Gamma}_{(N_{i})}$. If $(N_{i})$ is the collection of all such subgroups, we just write $\widehat{\Gamma}$. If the $N_{i}$ are normal subgroups, there is a group structure on $\widehat{\Gamma}_{(N_{i})}$ and an action $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$ given by the quotient map and left multiplication. Otherwise, there is still an action $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$ by left multiplication on left cosets in each coordinate. The canonical quotient map $p_{i}:\widehat{\Gamma}_{(N_{i})}\to\Gamma/N_{i}$ is continuous for the profinite topology and equivariant. The profinite topology can also be thought of as the subspace topology inherited from the Tychonoff topology on $\prod_{i}\Gamma/N_{i}$. Notice that $\widehat{\Gamma}_{(N_{i})}$ is closed in $\prod_{i}\Gamma/N_{i}$. In fact, if $\phi:\Gamma\to\prod_{i}\Gamma/N_{i}$ is the map coming from all the quotient maps, $\widehat{\Gamma}_{(N_{i})}=\overline{\phi(\Gamma)}$ (using that $(N_{i})$ is directed).
3 Odometers and Box Spaces
As we have seen, asymptotic dimension and DAD are closely related at least when the space being acted on is discrete. This relationship is encapsulated by 2.16 and, together with some topological facts about profinite completions, can be used to show the DAD (according to the definition for free actions) of an odometer formed from a sequence $(N_{i})$ of subgroups is the same as the asymptotic dimension of the box space formed from $(N_{i})$. Consequently, we see that questions about asymptotic dimension of such box spaces are special cases of questions about DAD. This connection is spiritually similar to how the DAD of an action is bounded above by the asymptotic dimension of its warped cone (this can be shown directly, but see [warpedcones] for definitions and a more complete discussion of warped cones and DAD).
Theorem 3.1.
Let $\Gamma$ be a countable group and $(N_{i})$ ($i\in I$) a countable collection of finite-index subgroups of $\Gamma$ directed by inclusion. Then $\text{DAD}_{free}(\Gamma\curvearrowright\Gamma_{(N_{i})})\leq\text{asdim}\Box_{(N_{i})}\Gamma$.
Proof.
Suppose $\text{asdim}\Box_{(N_{i})}\Gamma\leq d$ and fix a finite subset $F\in\Gamma$. Find $R>0$ so that $d_{\Gamma}(\gamma\cdot x,x)<R$ for all $\gamma\in F$ and $x\in\Gamma$, hence $d_{\Gamma/N_{i}}(\gamma\cdot x,x)<R$ for all $\gamma\in F$ and all $i$. Suppose $\text{asdim}\Box_{(N_{i})}\Gamma\leq d$. Then $\text{asdim}(\Gamma/N_{i})\leq d$ uniformly. Let $\mathcal{U}^{i}=\{U^{i}_{0},\ldots,U^{i}_{n}\}$ be a $(d,R,M_{R})$-cover of $\Gamma/{N_{i}}$. Then by 2.16 each $\mathcal{U}^{i}$ is also a $(d,F,B_{e}^{M_{R}+1}(\Gamma/N_{i}))$-cover for $\Gamma\curvearrowright\Gamma/{N_{i}}$ (and since the cardinality of balls in finite quotients of $\Gamma$ is bounded by the cardinality of balls in $\Gamma$, this bound is independent of $i$). If $K=\ker(\Gamma\to\widehat{\Gamma}_{(N_{i})})$, then $\Gamma/K\curvearrowright\widehat{\Gamma}_{(N_{i})}$ freely and is orbit equivalent to the original $\Gamma$-odometer. Find $i_{F}$ so that the quotient $\Gamma/K\to\Gamma/N_{i_{F}}$ induces a map $C_{F}(\Gamma/K)\to C_{F}(\Gamma/N_{i_{F}})$ which is an isometry on balls of radius $\leq M_{R}+2$ (this is possible, for isntance, by [boxspacesDT, Proposition 2.1]). Then pulling back $\mathcal{U}^{i_{F}}$ by the projection $p_{i_{F}}:\widehat{\Gamma}_{(N_{i})}\to\Gamma/N_{i_{F}}$ gives a $(d,F,B_{e}^{M_{R}+1}(\Gamma/N_{i_{F}}))$-cover for $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$. ∎
It will be important for 3.3 that the collection $(N_{i})$ is a sequence, meaning the index set is the natural numbers and $N_{i+1}\subset N_{i}$.
Lemma 3.2.
Let $\Gamma$ be a countable group and $(N_{i})$ a countable collection of finite index subgroups which is directed by inclusion. A subset $U\subset\widehat{\Gamma}_{(N_{i})}$ is clopen iff it is of the form $p_{i}^{-1}(S)$ where $S\subset\Gamma/N_{i}$ is any subset.
Proof.
First, sets of the form $p_{i}^{-1}(S)$ are clopen as $S\subset\Gamma/N_{i}$ is clopen and $p_{i}$ is continuous.
Fix the metric $d((x_{i}),(y_{i}))=\sum_{i=1}^{\infty}\frac{d_{\Gamma/N_{i}}(x_{i},y_{i})}{2^{i}\text{diam}(\Gamma/N_{i})}$ on $\widehat{\Gamma}$. If $U$ is clopen, then $\widehat{\Gamma}=U\cup U^{c}$ where $U$ and $U^{c}$ are separate connected components. In fact, since $\widehat{\Gamma}$ is a compact metric space, there is $\epsilon$ such that the $\epsilon$-neighborhoods of $U$ and $U^{c}$ do not intersect.
Now, suppose $U$ is not of the desired form. Then for all $i$ there is $j\geq i$ and $x_{j}\in\widehat{\Gamma}$ such that $p_{j}(x_{j})\in p_{j}(U)$ but $p_{j+1}(x_{j})\notin p_{j+1}(U)$. But then $x_{j}\notin U$ for all such (inifinitely-many) $j$ and $d(x_{j},U)\leq\sum_{n=j+1}^{\infty}\frac{1}{2^{n}}\to 0$ as $j\to\infty$ [^1] Technically, the sum is over a rearrangement, but that doesn’t matter.. Hence, there are points in $U^{c}$ arbitrarily close to $U$. By what was said above, this implies $U$ is not clopen. ∎
Proposition 3.3.
Let $\Gamma$ be a countable group and $(N_{i})$ a sequence of finite-index subgroups which is directed by inclusion (so $i\in\mathbb{N}$ and $N_{i+1}\subset N_{i}$). Let $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$ be the odometer action induced by left multiplication on each finite quotient. If $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})\leq d$ then $\text{asdim}(\Box_{(N_{i})}\Gamma)\leq d$.
Proof.
Fix a proper, left-invariant metric on $\Gamma$. Let $R\geq 0$ be given and set $F=B_{e}^{R}(\Gamma)$. Let $\mathcal{V}=\{V_{0},\ldots,V_{d}\}$ be an open $(d,F,S)$-cover for $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$. Observe that for each $i$, the distance between the $F$-components of $p_{i}(V_{j})$ is >R>R according to the metric on $\Gamma/N_{i}$.
Now by 2.8 we can assume each $V_{j}$ is clopen and hence by 3.2 is the pullback of $T\subset\Gamma/N_{i}$ for some $i$. Find $i_{0}$ such that all of the $V_{j}$ are pullbacks of subsets of $\Gamma/N_{i_{0}}$. Then if $i\geq i_{0}$, $p_{i}(x)\in p_{i}(V_{j})$, and $f\cdot p_{i}(x)\in p_{i}(V_{j})$, then $x,f\cdot x\in V_{j}$. Thus, each $F$-component of $p_{i}(V_{j})$ is $S$-bounded for all $i\geq i_{0}$; so by 2.16 the diameter (as a subset of $\Box_{(N_{i})}\Gamma$) of each $R$-component of $p_{i}(V_{j})\subset\Gamma/N_{i}$ is bounded by $\text{diam}(S)$ for $i\geq i_{0}$. This gives a $(d,R,\text{diam}(S))$-cover of each $\Gamma/N_{i}$ for all $i\geq i_{0}$. Since $(N_{i})$ is a sequence, the set $\{i\mid i<i_{0}\}$ is finite, and so the diameter of $\Gamma/N_{i}$ for $i<i_{0}$ is uniformly bounded. We have therefore shown $\text{asdim}(\Gamma/N_{i})\leq d$ uniformly. The last part then follows from 2.14. ∎
We now state both inequalities (when $(N_{i})$ is a sequence) in one theorem:
Theorem 3.4.
Let $\Gamma$ be a countable group and $(N_{i})$ a sequence of finite index subgroups which is directed by inclusion. Suppose $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$ by left multiplication. Then $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\text{asdim}\Box_{(N_{i})}\Gamma$. ∎
Corollary 3.5.
Let $\Gamma$ be a countable, residually finite group and $(N_{i})$ a sequence of finite index subgroups which is also a filtration of $\Gamma$ (see 2.11). Then $\text{DAD}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\text{asdim}\Box_{(N_{i})}\Gamma$. ∎
This can also be combined with [boxspacesDT, Proposition 3.1]:
Corollary 3.6.
Let $\Gamma$ be a finitely generated, residually finite group and $(N_{i})$ a sequence of finite index subgroups which is also a filtration of $\Gamma$. Then either $\text{DAD}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\infty$ or $\text{DAD}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\text{asdim}\Gamma$. The dimension is finite if $\Gamma$ has polynomial growth (equivalently is virtually nilpotent). ∎
4 Product Theorem
Now we begin our study of permanence properties of DAD which will lead to our main result. Most basically, one would expect DAD to be subadditive over products of groups acting on products of spaces, and the proof of this fact will exhibit some of the same ideas used in the next, more technical, section. A product theorem for the dynamic asymptotic dimension of groupoids was recently established by Bönicke in [bonicke2023dynamic], but we will go on to prove the more general Hurewicz mapping theorem for the transformation group case. It is worth remarking that the product theorem for groupoids does not really require any additional ideas from the case for group actions.
We rely heavily on the ideas of [Brodskiy2006AHT] for both this section and the next and have noted when one of our lemmas is inspired by one of theirs. Indeed, if one relaxes the definition of DAD to allow Borel sets, or even sets obtained from finite operations with open sets, the work of [Brodskiy2006AHT] can be transported to the dynamical setting with very little extra work. Our solution to the extra technicality created by the topology of $X$ comes from a lemma provided in this section which we refer to as ‘the trick’. This lemma capitalizes on some rigidity to provide a way of switching between covers using open or closed sets while preserving the properties related to DAD. Part (2) was previously established in [toprigiditydad, Lemma 2.8(i)], and the essential content of part (1) also appeared in the surrounding discussion in that paper.
Definition 4.1.
Let $\Gamma\curvearrowright X$ be an action on a compact space by homeomorphisms and $A\subseteq X$. If $k\geq n+1\geq 1$, an $(n,k)$-dimensional control function for $A$ (with respect to $\Gamma\curvearrowright X$) is a function $\mathcal{D}_{A}:\mathcal{P}_{\text{fs}}(\Gamma)\to\mathcal{P}_{\text{fs}}(\Gamma)$ such that for any $F\in\mathcal{P}_{\text{fs}}(\Gamma)$ there is a $(k-1,F,\mathcal{D}_{A}(F))$-cover $\mathcal{U}=\{U_{0},\ldots,U_{k-1}\}$ of $A$ (with respect to $\Gamma\curvearrowright X$) by sets open in the subspace topology on $A$ such that each $x\in A$ belongs to at least knk-n elements of $\mathcal{U}$ (equivalently $\cup_{i\in T}U_{i}=A$ for every $T\subset\{0,\ldots,k-1\}$ consisting of $n+1$ elements).
Lemma 4.2.
(the trick) Let $\Gamma\curvearrowright X$ be an action on a compact metric space by homeomorphisms and $A\subset X$ closed.
  • (1)
    Suppose $F,S\in\mathcal{P}_{\text{fs}}(\Gamma)$, $n\geq 0$, $k\geq n+1$, and $\mathcal{U}=\{U_{0},\ldots,U_{k-1}\}$ is a (k1,F,S)(k-1,F,S)-cover for $A$ by sets open in the subspace topology on $A$ such that each $x$ belongs to $\geq k-n$ elements. Then there is a (k1,F,S)(k-1,F,S)-cover by sets which are the closures of open sets (so closed in $X$) and still have the property that each $x$ belongs to $\geq k-n$ elements.
  • (2)
    Suppose the $F$-components of $A$ are $S$-bounded. Then $A$ is contained in a set which is open in $X$ and whose $F$-components are $S$-bounded.
Proof.
For the first part, assume $k>1$ and let $\mathcal{U}=\{U_{0},\ldots,U_{k-1}\}$ be such a cover. Then $\bigcup_{S\subset\{0,\ldots,k\}:|S|=k-n}\cap_{s\in S}U_{s}$ is an open cover, and so has some Lebesgue number $\lambda$. This shows every $x\in X$ has a $\lambda$-ball which is contained in at least knk-n elements of $\mathcal{U}$. Replace the $U_{i}$ with the closures of their $\lambda/2$-interiors. If $n=0$ and $k=1$, $\mathcal{U}=\{A\}$, which is already a cover by closed sets (closed in $X$ in fact).
We can assume $S\in\mathcal{P}_{fs}(\Gamma)$. To prove $(2)$, suppose not, that is, suppose for every $n$, the $F$ components of $N_{1/n}(U)$ are not $S$-bounded. So for every $n$, we have an $F$-chain $x^{(n)}_{0},\ldots,x^{(n)}_{K_{n}}$ in $N_{1/n}(U)$ with $x^{(n)}_{0},\ldots,x^{(n)}_{K_{n}-1}\in S\cdot x_{0}$ but $x_{K_{n}}=\gamma_{n}\cdot x^{(n)}_{0}$ for $\gamma_{n}=f_{n}s_{n}$ with $s_{n}\in S$, $f_{n}\in F$, and $\gamma_{n}\notin S$. Since there are finitely-many $\gamma\in\Gamma$ with the property that $\gamma=fs$ for $s\in S$ and $f\in F$, we must have $\gamma_{n}=\lambda$ for infinitely-many $n$. Furthermore, there are finitely-many sequences $f_{1},\ldots,f_{K_{n}}$ ($f_{k}\in F$) such that $f_{k}\cdots f_{1}\in S$ for all $1\leq k<K_{n}$ but $f_{K_{n}}\cdots f_{1}=\lambda\notin S$, and so for infinitely-many $n$ we have an $F$-chain $x^{(n)}_{0},\ldots,x^{(n)}_{K}$ in $N_{1/n}(U)$ and $f_{1},\ldots,f_{K}\in F$ such that $f_{k+1}\cdot x^{(n)}_{k}=x^{(n)}_{k+1}$ for $0\leq k<K$, and also $x_{K}=\lambda\cdot x_{0}$ with $\lambda\notin S$. Since $X$ is compact, we have a subsequence $x_{0}^{(n_{j})}$ converging to a point $y_{0}\in X$, and so $f_{k}\ldots f_{1}\cdot x_{0}^{(n_{j})}=x_{k}^{(n_{j})}$ converges to $f_{k}\ldots f_{1}\cdot y_{0}$ for all $1\leq k\leq K$ by continuity. Since $x_{k}^{(n_{j})}\in N_{1/n_{j}}(U)$ and $U$ is closed in $X$, $x_{k}^{(n_{j})}$ converges to a point in $U$. Hence, we have an $F$-chain in $U$ which is not $S$-bounded. ∎
Remark 4.3.
We will frequently apply the lemma above by simply saying ‘use the trick’, when we want to replace an open cover by a closed one or a closed set with a larger open set. For a closed subspace $A\subset X$, notice that the trick implies it is equivalent to define $\text{DAD}(A)$ and related terms using covers by sets which are open in $X$ (the definition in 2.6, used sets which are open in the subspace topology on $A$).
Lemma 4.4.
(compare to [Brodskiy2006AHT, Theorem 2.4]) Suppose $A\subset X$ is closed. Suppose $\mathcal{D}_{A}^{(n+1)}$ is an $n$-dimensional control function for $A$. Define $\{\mathcal{D}_{A}^{(i)}\}_{i\geq n+1}$ inductively on powers of $F$ by $\mathcal{D}_{A}^{(i+1)}(F^{r}):=F^{r}\cdot\mathcal{D}^{(i)}_{A}(F^{3r})\cdot F^{r}$; then for each $k$, there is an open $(k-1,F^{r},\mathcal{D}^{(k)}_{A}(F^{r}))$-cover of $A$ such that each $x\in A$ is covered by at least knk-n elements of the cover. (We say it this way since the functions $\mathcal{D}_{A}^{(k)}$ are not defined on all elements of $\mathcal{P}_{\text{fs}}(\Gamma)$ unless $\Gamma$ is finitely generated, so we can’t quite call them control functions).
Proof.
The case $k=n+1$ is obvious. Suppose the result holds for some larger $k$. Let $\mathcal{U}=\{U_{1},\ldots,U_{k}\}$ be an open $(k-1,F^{3r},\mathcal{D}_{A}^{(k)}(F^{3r}))$-cover for $A$ with all $x\in A$ in at least knk-n elements of $\mathcal{U}$. Using (1) from the trick (see 4.3), we can replace the sets $U_{i}$ by the closures of open sets, $V_{i}$, which form a cover of $A$ with the same properties. Define $U_{i}^{\prime}=(F^{r}\cdot U_{i})\cap A$ (these are closed sets). Then $\mathcal{U}^{\prime}=\{U_{1}^{\prime},\ldots,U_{k}^{\prime}\}$ is a closed $(k-1,F^{r},F^{r}\cdot\mathcal{D}_{A}^{(k)}(F^{3r})\cdot F^{r})$-cover for $A$. Notice that $F^{r}\cdot\mathcal{D}_{A}^{(k)}(F^{3r})\cdot F^{r}=\mathcal{D}_{A}^{(k+1)}(F^{r})$.
Define $U_{k+1}^{\prime}$ to be the union of all the (disjoint, open) sets $W_{S}:=(\cap_{s\in S}V_{s})\setminus\cup_{i\notin S}U_{i}^{\prime}$ where $S\subset\{1,\ldots,k\}$ has knk-n elements. We claim that for $S\neq T$ (where $S,T\subset\{1,\ldots,k\}$ each have knk-n elements), the sets $W_{S}$ and $W_{T}$ are $F^{r}$-disjoint (i.e. not connected by an $F^{r}$-chain contained in their union).
Suppose $a,b\in U_{k+1}^{\prime}\subset A$ are such that $a\in\cap_{t\in T}V_{t}\setminus\cup_{i\notin T}U_{i}^{\prime}$ and $b\in\cap_{s\in S}V_{s}\setminus\cup_{i\notin S}U_{i}^{\prime}$ for $T\neq S$ and |S|=|T|=kn|S|=|T|=k-n. Then there is $t\in T\setminus S$ such that $a\in V_{t}$, so if $f\cdot a=b$ for some $f\in F^{r}$, then $b\in U_{t}^{\prime}\supseteq F^{r}\cdot V_{t}$, a contradiction.
Now suppose $x\in A$ belongs to exactly knk-n sets $U_{i}^{\prime}$, $i\leq k$. Since the sets $\cap_{s\in S}V_{s}$ form a cover as $S$ ranges over subsets of $\{1,\ldots,k\}$ with |S|=kn|S|=k-n; $x\in\cap_{s\in S}V_{s}$ for some $s\in S$. But then if $x\notin U_{k+1}^{\prime}$, we must have $x\in U^{\prime}_{j}$ for some $j\notin S$, contradicting that $x$ is only contained in knk-n sets $U_{i}^{\prime}$.
Thus, $\{U^{\prime}_{k+1},U^{\prime}_{k},\ldots,U^{\prime}_{1}\}$ is a $(k,F^{r},\mathcal{D}_{A}^{(k+1)}(F^{r}))$-cover with each $x\in A$ contained in at least kn+1k-n+1 sets (recall that $U_{k+1}^{\prime}$ is open and the other sets are closed). We can then apply the trick (2) to each closed set to produce a cover by open sets with the same properties. ∎
Theorem 4.5.
(compare to [Brodskiy2006AHT, Theorem 2.5]) If $\Gamma\curvearrowright X$ and $\Lambda\curvearrowright Y$ are free, then $\text{DAD}(\Gamma\times\Lambda\curvearrowright X\times Y)\leq\text{DAD}(\Gamma\curvearrowright X)+\text{DAD}(\Lambda\curvearrowright Y)$.
Proof.
Fix $F\in\mathcal{P}_{\text{fs}}(\Gamma)$ and $H\in\mathcal{P}_{\text{fs}}(\Lambda)$. Suppose $\Gamma\curvearrowright X$ and $\Lambda\curvearrowright Y$ have dimension at most $m$ and at most $n$, respectively, and let $\mathcal{D}_{X}$ and $\mathcal{D}_{Y}$ be respective control functions. Let $k=m+n+1$. Use 4.4 to produce the functions $\mathcal{D}^{(k)}_{X}$ and $\mathcal{D}^{(k)}_{Y}$. So there is an open $(k-1,F,\mathcal{D}^{(k)}_{\Gamma\curvearrowright X}(F))$-cover $\mathcal{U}=\{U_{i}\}_{i=0}^{k-1}$ of $X$ and an open $(k-1,H,\mathcal{D}^{(k)}_{\Lambda\curvearrowright Y}(H))$-cover $\mathcal{V}=\{V_{i}\}_{i=0}^{k-1}$ of $Y$ such that each $x\in X$ is in at least kmk-m elements of $\mathcal{U}$ and each $y\in Y$ is in at least knk-n elements of $\mathcal{V}$. Since km+kn=k+1k-m+k-n=k+1, the family $\{U_{i}\times V_{i}\}_{i=0}^{k-1}$ is a cover of $X\times Y$. Moreover, it is a $(F\times H,\mathcal{D}^{(k)}_{\Gamma\curvearrowright X}(F)\cdot\mathcal{D}^{(k)}_{\Lambda\curvearrowright Y}(H))$-cover for $\Gamma\times\Lambda\curvearrowright X\times Y$. ∎
5 Hurewicz-type Theorem
In this section, we will prove a more general permanence property witnessing a natural dimension-theoretic property of DAD. We start by defining the dimension of an appropriate morphism in the category of group actions. As dynamic asymptotic dimension is meant to measure the large scale complexity of orbits, the appropriate kind of morphism here is one which sends orbits to orbits and induces a coarse map of the orbit structures. The following definition is not new, but it was inspired by the definition of continuous orbit couple in [LiDynamics, Definition 2.6]. A continuous orbit couple induces a ‘coarse equivalence of orbits’, so we weaken this definition to one which will give a ‘coarse map’ of orbits.
Definition 5.1.
Suppose $\Gamma\curvearrowright X$ and $\Lambda\curvearrowright Y$ are actions on compact metric spaces by homeomorphisms and that $f:X\to Y$ and $\alpha:\Gamma\times X\to\Lambda$ are continuous maps with $f(\gamma\cdot x)=\alpha(\gamma,x)\cdot x$. Define $\text{DAD}(f)=\sup_{A\subset Y:\text{DAD}(A)=0}\text{DAD}_{free}(f^{-1}(A))$ (with $\text{DAD}(f^{-1}(A))$ defined as in 2.6).
The intuition behind this definition is we want to measure the largest dimension of a subspace of $X$ which $f$ collapses down to something zero dimensional in the image. Recall for instance that in the category of vector spaces, this would be the nullity of a linear map, i.e. the dimension of the pullback of $\{0\}$. In the category of topological spaces, one can similarly define the dimension of a function to be the supremum over the dimensions of pullbacks of points. However, as in the case of asymptotic dimension, we must alter this definition slightly by considering pullbacks of arbitrary 0-dimensional subspaces. See [Brodskiy2006AHT, Proposition 4.1] for an example which shows this deviation is necessary.
Many of the ideas in this section come from [Brodskiy2006AHT], so as in the previous section we note at the beginning of each statement what lemma is being analogized.